# Nonlinear skin breathing modes in one-dimensional nonreciprocal mechanical lattices

### v1

@@ Abstract
We investigate the interplay of nonreciprocity and nonlinearity in a one-dimensional nonlinear Klein–Gordon chain of classical oscillators coupled by asymmetric springs, akin to a mechanical analogue of the Hatano–Nelson model with onsite nonlinearity. Using multiple-scale analysis, we show that families of nonlinear skin breathing modes—time-periodic, boundary-localized oscillations—emerge from their linear counterparts at any nonreciprocal strength. We derive an explicit nonlinear frequency shift for these families of nonlinear breathing modes, showing its dependence on amplitude, nonlinearity type, lattice size, and nonreciprocity, and we predict the emergence of genuine skin end breathers at the boundary once their oscillation frequency and higher harmonics enter the spectral gaps of the linear spectrum. Numerical pseudo-arclength continuation confirms full families of solutions for both hardening and softening nonlinearities. Furthermore, the Floquet analysis shows that these modes can be either linearly stable or unstable, with Floquet eigenvectors exhibiting skin localization inherited from the asymmetric couplings. Our results extend the nonlinear non-Hermitian skin effect from stationary modes to intrinsically time-periodic excitations, providing a pathway to engineer and control breathing modes in nonreciprocal mechanical metamaterials.
@@ I Introduction
Nonreciprocal mechanical metamaterials have recently attracted significant attention [1, 2, 3, 4, 5, 6, 7, 8]. These metamaterials provide a versatile platform to explore non-Hermitian physics and wave control, as they can in general be effectively modeled by simple mass–spring systems in which the left- and right-going spring couplings differ from one another. Among the most fundamental features of these systems is the non-Hermitian skin effect (NHSE), whereby all of the normal modes collapse at one edge of the system under open boundary conditions (OBCs) [9, 10, 11, 12, 13, 14, 15]. The NHSE was first demonstrated in experiment in mechanical systems [1, 16] before finding analogues in electronics [17, 18], acoustics [19, 20], optics [21, 22], and other phononic settings [23]. This provide a compelling motivation for a deeper investigation of nonreciprocal mechanical systems.
Nevertheless, research on the interplay between the NHSE and nonlinearity has largely overlooked mechanical systems, focusing instead on photonic lattices like the Hatano–Nelson (HN) model with onsite nonlinearity [24, 25, 26, 27, 28, 29, 30]. Through these studies it was found that nonlinear NHSE persists for finite-amplitude waves. More specifically, perturbation methods clarified in theory and numerical simulations the emergence of nonlinear skin modes from their linear counterparts as the nonlinearity grows [26, 27, 29]. However, the particularity of these modes is that they are stationary, a property inherent to systems governed by first-order differential equations such as $\dot{y}=Hy$. Indeed, solutions in such systems are of the form $y(t)=Ue^{-i\Omega t}$ reducing their dynamics to the eigenvalue problem $Hu=\Omega u$. It follows that, the field profile $\lvert y(t)\rvert=U$ becomes time-independent [26, 27, 29].
@@
In mechanical lattices, however, the dynamics is governed by second-order equations like, $\ddot{y}=Hy$. It follows that when looking for solutions of the form, $y(t)=Ue^{-i\Omega t}$ yields $\ddot{y}=-\Omega^{2}y$ with $HU=-\Omega^{2}U$. Thus, the system dynamics necessarily contains $e^{\pm i\Omega}$ terms leading to a general form of the solution with two components $y(t)=U^{\prime}\cos(\Omega t)+U^{\prime\prime}\sin(\Omega t)$, whose amplitudes $\lvert y(t)\rvert$ oscillate in time. It follows that the nonlinear modes in mechanical systems are inherently time-periodic (like discrete breathers [31, 32, 33, 34]) rather than stationary. This difference raises several fundamental questions. Do nonlinear skin breathing (oscillating) modes exist in nonreciprocal mechanical lattices? If so, how do their frequencies and envelopes depend on control parameters like the nonlinearity, nonreciprocity, and lattice size?
We address these questions by studying a one-dimensional (1D) chain of nonlinear classical oscillators connected through asymmetric spring couplings, referred to as the nonlinear nonreciprocal Klein–Gordon (KG) chain. This model originates from the mechanical analogue of the celebrated Hatano–Nelson (HN) chain [35, 36] with a cubic onsite nonlinearity [29]. Mechanical implementations of closely related systems already exist, including nonreciprocal variants of the discrete sine–Gordon model [5, 8] (see also [18]). Notably, Ref. [8] investigates, in the presence of dissipation, traveling breathers within the bulk of the lattice. By contrast, the present work focuses on nonlinear normal modes that originate from the linear skin modes. It follows that by definition, these states are standing, time-periodic, and localized at the edge selected by the NHSE.
@@ II Nonlinear skin breathing modes
Figure 3: (a) Dependence of the nonlinear frequency $\Omega$ on the amplitude $\lVert\vec{y}\rVert^{2}$ for families of nonlinear skin breathing modes emerging from the linear skin modes with $j=1$ (red), $j=6$ (blue), $j=10$ (cyan), $j=14$ (yellow), $j=18$ (purple) and $j=24$ (green), shown from the bottom to the top curves. The dashed portions of the curves indicate unstable modes, while the solid portions correspond to stable ones (this representation is qualitative). The parameters used in the computations are $N=24$, $\alpha=1$, $\gamma=0.25$, and $g=+1$ (hardening). The numerical continuation is stopped when $\lVert\vec{y}\rVert^{2}=3$, which roughly corresponds to weak and moderate nonlinear strengths. (b) Profile of a representative nonlinear skin breathing mode with $\lVert\vec{y}\rVert^{2}=0.2$, belonging to the family with $j=24$, see star symbol. Note that its linear counterpart is shown in the upper inset of Fig. 2(a). (c) Time evolution of the state in (b) used as the initial condition, clearly demonstrating an amplitude that oscillates in time. (d) Same as in (a), but for g=1g=-1 (softening). (e) Same as in (b), but for the softening case shown in (d), see dot symbol. Note that its linear counterpart is depicted in the lower inset of Fig. 2(a). (f) Same as in (c), but using the state in (e) as the initial condition.
@@
Furthermore, as the dynamical matrix is non-Hermitian [Appendix A.1], we can also define a left eigenvalue problem $\omega^{2}\vec{v}=D^{T}\vec{v}$, with $\vec{v}=(v_{1},v_{2},\ldots,v_{N-1},v_{N})^{T}$. The right and left eigenvectors associated with a frequency $\omega_{j}$,
$u_{j,n}=r^{\frac{n}{2}}\sin\left(\frac{nj\pi}{N+1}\right),\quad v_{j,n}=r^{-\frac{n}{2}}\sin\left(\frac{nj\pi}{N+1}\right)$ (4)
@@
We now address whether families of nonlinear normal modes satisfying Eq. (1) can emerge from the linear skin modes as their amplitudes increase. A schematic illustration of this procedure and the expected outcome is shown in Fig. 2(b), where the parameter $\gamma$ is fixed while the frequency, $\Omega$ is treated as a free parameter. We apply the method of multiple scales, seeking solutions of the form [37, Chap. 6]
@@
where $\epsilon\ll 1$ and the $\{^{\star}\}$ stands for the complex conjugate operation. In addition, the $t$ and $T=\epsilon^{2}t$ are fast and slow time variables. We find that at order $\epsilon$, $\Omega^{2}w_{n}=-\sum_{m}D_{n,m}w_{m}$. Consequently, $\Omega=\omega_{j}$ and $w_{n}=u_{j,n}$, indicating that the ansatz correctly reproduces the linear frequencies and the associated skin modes in the small amplitude limit. Proceeding to order $\epsilon^{3}$, the condition for no resonant terms leads to the amplitude equation $2i\omega_{j}\partial_{T}A+3\mu\lvert A\rvert^{2}A=0$, with $\mu=2g\left(N+1\right)^{-1}\left(\sum_{n=1}^{N}v_{n}u_{n}^{3}\right)$. It follows that,
@@
Let us now express the $\epsilon$ in terms of the macroscopic parameter controlling the nonlinearity in our system, like the total amplitude $\lVert\vec{y}\rVert^{2}$. Indeed, the $\lVert\vec{y}\rVert^{2}$ is a constant of motion along the periodic or quasi-periodic orbit defined by the nonlinear breathing mode, $y_{n}(t)$. It follows that Eq. (5), leads to $4\epsilon^{2}\lvert A_{0}\rvert^{2}=\lVert\vec{y}\rVert^{2}/\lVert\vec{u}\rVert^{2}$ when considering only the first order in $\epsilon$. Consequently,
@@
Lets us now perform numerical verifications of the theory above. We implement a pseudo-arclength solver [38, 39, 40] to numerically find families of exact nonlinear solutions of Eq. (1). The initial point for the pseudo–arclength solver is obtained using a shooting method, which searches for a small–amplitude exact solution with frequency $\Omega$, using the corresponding linear skin mode as an initial guess [39, 40]. For the direct numerical simulations of the equations of motion, we use a Runge–Kutta scheme of order $8$ based on the Dormand–Prince method (DOP853853), with a time step $\Delta t=0.05$ and a one-step tolerance of $10^{-11}$ [41, 42, 43]. Hereafter, we consider $N=24$, $\alpha=1$, $\gamma=0.25$ and terminate the pseudo-arclength continuation at $\lVert\vec{y}\rVert^{2}=3$ so as to avoid intricate bifurcation structures that emerge at large nonlinearities, which are beyond the scope of this work.
Figure 4: Dependence of the nonlinear frequency sensitivity factor $\mathcal{S}=\partial\widetilde{\Omega}/\partial S$ on the lattice size NN, where $\widetilde{\Omega}=(\Omega-\omega)/\omega$ and $S=\lVert\vec{y}\rVert^{2}$. We compute $\mathcal{S}$ for fixed values $\gamma=0.25$ (red squares), $\gamma=0.4$ (blue dots), $\gamma=0.55$ (cyan triangles), and $\gamma=0.7$ (purple stars). The derivatives are evaluated using a two-point finite difference at $\lVert\vec{y}\rVert^{2}=0.01$ and 0 [44]. The dashed line is a guide to the eye showing $\mathcal{S}(N)\sim\exp(-2.5\,N)$ (see text for details). Inset: Dependence of $\mathcal{S}$ on $\gamma$ for three lattice sizes, $N=24$, $N=32$, and $N=42$, shown as the blue, red, and cyan dotted symbols, respectively. We find that $\mathcal{S}(\gamma)\sim r^{-2}$, $r^{-2.5}$ and $r^{-3}$ for $N=24$, $N=32$, and $N=42$ respectively. In all cases, the analytical predictions show as solid curves overlapping with the numerical results, see also Fig. 7.
@@
We numerically compute the dependence of this frequency sensitivity factor, $\mathcal{S}$ [Eq. (9)] against the lattice size, $N$ in case of the hardening nonlinearitly with $g=1$. Note that the first derivative is evaluated using a two points finite difference formula in the interval $\lVert\vec{y}\rVert^{2}\in(0,0.01]$ [44]. Figure 4 shows the results of these computations for the family of nonlinear skin breathing modes with $j=1$ considering four different values of the nonreciprocity, $\gamma=0.25$, $\gamma=0.4$, $\gamma=0.55$, and $\gamma=0.7$, corresponding respectively to the red squares, blue dots, cyan triangles, and purple stars in Fig. 4. We find that for all values of $\gamma$, the $\mathcal{S}$ tends to asymptotically decay as the lattice size growth. The $\mathcal{S}\sim\exp\left(q\,N\right)$ with $q=q(\gamma)$. Interestingly, for $\gamma=0.4,~0.55,~0.7$, we find that the power exponent $q(\gamma)$ tends to saturate at $q\approx-2.5$, as indicated by the black dotted line in Fig. 4. Note that this exponent is obtained by fitting the last eight points of the numerical data.
Moreover, the inset of Fig. 4 depicts similar computations as above, but varying this time the nonreciprocal strength $\gamma$ at fixed lattice sizes, $N=24$, $N=32$ and $N=42$ respectively the red squares, blue diamond and cyan triangles. Following similar arguments as above, we expect $\mathcal{S}\sim r^{p}$, with $p=p(N)$. By numerical fitting the sensitivity using the power law $\mathcal{S}=\sigma r^{-p}$, we find that $\mathcal{S}(\gamma)\sim r^{-2}$, $r^{-2.5}$, and $r^{-3}$ for $N=24$, $N=32$, and $N=42$, respectively [46]. The analytical predictions from Eq. (9) (solid curves) are superimposed on the numerical data and show excellent agreement, consistent with Fig. 7 in Appendix B.

### v2

@@ Abstract
We investigate the interplay of nonreciprocity and nonlinearity in a one-dimensional nonlinear Klein-Gordon chain of classical oscillators coupled by asymmetric springs, akin to a mechanical analogue of the Hatano–Nelson model with onsite nonlinearity. Using multiple-scale analysis, we show that families of nonlinear skin breathing modes—time-periodic, boundary-localized oscillations—emerge from their linear counterparts at any nonreciprocal strength. We derive an explicit nonlinear frequency shift for these families of nonlinear breathing modes, showing its dependence on amplitude, nonlinearity type, lattice size, and nonreciprocity, and we predict the emergence of genuine skin end breathers at the boundary once their oscillation frequency and higher harmonics enter the spectral gaps of the linear spectrum. Numerical pseudo-arclength continuation confirms full families of solutions for both hardening and softening nonlinearities. Furthermore, the Floquet analysis shows that these modes can be either linearly stable or unstable, with Floquet eigenvectors exhibiting skin localization inherited from the asymmetric couplings. Our results extend the nonlinear non-Hermitian skin effect from stationary modes to intrinsically time-periodic excitations, providing a pathway to engineer and control breathing modes in nonreciprocal mechanical metamaterials.
@@ I Introduction
Nonreciprocal mechanical metamaterials have recently attracted significant attention [1, 2, 3, 4, 5, 6, 7, 8, 9]. These metamaterials provide a versatile platform to explore non-Hermitian physics and wave control, as they can in general be effectively modeled by simple mass–spring systems in which the left- and right-going spring couplings differ from one another. Among the most fundamental features of these systems is the non-Hermitian skin effect (NHSE), whereby all the normal modes collapse at one edge of the system under open boundary conditions (OBCs) [10, 11, 12, 13, 14, 15, 16]. The NHSE was first demonstrated in experiment in mechanical systems [1, 17] before finding analogues in electronics [18, 19], acoustics [20, 21], optics [22, 23], and other phononic settings [24]. These developments provide strong motivation for a deeper and systematic investigation of nonreciprocal mechanical systems.
Nevertheless, research on the interplay between the NHSE and nonlinearity has largely overlooked mechanical systems, focusing instead on photonic lattices like the Hatano–Nelson (HN) model with onsite nonlinearity [25, 26, 27, 28, 29, 30, 31]. Through these studies it was found that nonlinear NHSE persists for finite-amplitude waves. More specifically, perturbation methods clarified in theory and numerical simulations the emergence of nonlinear skin modes from their linear counterparts as the nonlinearity grows [27, 28, 30]. However, the particularity of these modes is that they are stationary, a property inherent from systems governed by first-order differential equations such as $i\dot{y}=Hy$. Indeed, solutions in such systems are of the form $y(t)=Ue^{-i\Omega t}$ reducing their dynamics to the eigenvalue problem $Hu=\Omega u$. It follows that, the field profile $\lvert y(t)\rvert=U$ becomes time-independent [27, 28, 30].
@@
In mechanical lattices, however, the dynamics is governed by second-order equations like, $\ddot{y}=Hy$. It follows that when looking for solutions of the form, $y(t)=Ue^{-i\Omega t}$ yields $\ddot{y}=-\Omega^{2}y$ with $HU=-\Omega^{2}U$. Thus, the system dynamics necessarily contains $e^{\pm i\Omega t}$ terms leading to a general form of the solution with two components $y(t)=U^{\prime}\cos(\Omega t)+U^{\prime\prime}\sin(\Omega t)$, whose amplitudes $\lvert y(t)\rvert$ oscillate in time. It follows that the nonlinear modes in mechanical systems are inherently time-periodic (like discrete breathers [32, 33, 34, 35]) rather than stationary. This difference raises several fundamental questions. Do nonlinear skin breathing (oscillating) modes exist in nonreciprocal mechanical lattices? If so, how do their frequencies and envelopes depend on control parameters like the nonlinearity, nonreciprocity, and lattice size?
We address these questions by studying a one-dimensional (1D) chain of nonlinear classical oscillators connected through asymmetric spring couplings, referred to as the nonlinear nonreciprocal Klein-Gordon (KG) chain. This model originates from the mechanical analogue of the celebrated Hatano–Nelson (HN) chain [36, 37] with a cubic onsite nonlinearity [30]. Mechanical implementations of closely related systems already exist, including nonreciprocal variants of the discrete sine–Gordon model [6, 9] (see also [19]). Notably, Ref. [9] investigates, in the presence of dissipation, traveling breathers within the bulk of the lattice. By contrast, the present work focuses on nonlinear normal modes that originate from the linear skin modes. It follows that by definition, these states are standing, time-periodic, and localized at the edge selected by the NHSE.
@@ II Nonlinear skin breathing modes
Figure 3: Dependence of the nonlinear frequency $\Omega$ on the amplitude $\lVert\vec{y}\rVert^{2}$ for families of nonlinear skin breathing modes emerging from the linear skin modes with $j=1$ (red), $j=6$ (blue), $j=10$ (cyan), $j=14$ (yellow), $j=18$ (purple) and $j=24$ (green), shown from the bottom to the top curves. The dashed portions of the curves indicate unstable modes, while the solid portions correspond to stable ones (this representation is qualitative). The parameters used in the computations are $N=24$, $\alpha=1$, $\gamma=0.25$, and $g=+1$ (hardening). The numerical continuation is stopped when $\lVert\vec{y}\rVert^{2}=3$, which roughly corresponds to weak and moderate nonlinear strengths. (b) Profile of a representative nonlinear skin breathing mode with $\lVert\vec{y}\rVert^{2}=0.2$, belonging to the family with $j=24$, see star symbol. Note that its linear counterpart is shown in the upper inset of Fig. 2(a). (c) Time evolution of the state in (b) used as the initial condition, clearly demonstrating an amplitude that oscillates in time. (d) Same as in (a), but for g=1g=-1 (softening). (e) Same as in (b), but for the softening case shown in (d), see dot symbol. Note that its linear counterpart is depicted in the lower inset of Fig. 2(a). (f) Same as in (c), but using the state in (e) as the initial condition.
@@
Furthermore, as the dynamical matrix is non-Hermitian (Appendix A.1), we can also define a left eigenvalue problem $\omega^{2}\vec{v}=D^{T}\vec{v}$, with $\vec{v}=(v_{1},v_{2},\ldots,v_{N-1},v_{N})^{T}$. The right and left eigenvectors associated with a frequency $\omega_{j}$,
$u_{j,n}=r^{\frac{n}{2}}\sin\left(\frac{nj\pi}{N+1}\right),~~v_{j,n}=r^{-\frac{n}{2}}\sin\left(\frac{nj\pi}{N+1}\right)$ (4)
@@
We now address whether families of nonlinear normal modes satisfying Eq. (1) can emerge from the linear skin modes as their amplitudes increase. A schematic illustration of this procedure and the expected outcome is shown in Fig. 2(b), where the parameter $\gamma$ is fixed while the frequency, $\Omega$ is treated as a free parameter. We apply the method of multiple scales, seeking solutions of the form [38, Chap. 6]
@@
where $\epsilon\ll 1$ and the $\{^{\star}\}$ stands for the complex conjugate operation. In addition, the $t$ and $T=\epsilon^{2}t$ are fast and slow time variables. We find that at order $\epsilon$, $\Omega^{2}w_{n}=\sum_{m}D_{n,m}w_{m}$. Consequently, $\Omega=\omega_{j}$ and $w_{n}=u_{j,n}$, indicating that the ansatz correctly reproduces the linear frequencies and the associated skin modes in the small amplitude limit. Proceeding to order $\epsilon^{3}$, the condition for no resonant terms leads to the amplitude equation
$2i\omega_{j}\frac{\partial A}{\partial T}+3\mu\lvert A\rvert^{2}A=0,$ (6)
with $\mu=2g\left(N+1\right)^{-1}\left(\sum_{n=1}^{N}v_{j,n}u_{j,n}^{3}\right)$. It follows that,
@@
Let us now express $\epsilon$ in terms of a macroscopic parameter controlling the nonlinearity in our system, namely the squared maximum displacement, $\lVert\vec{y}\rVert^{2}$ (hereafter referred to as the total amplitude), which is a constant coefficient. Indeed, we see that the complex amplitude equation [Eq. (6)] is a Hermitian nonlinear Schrödinger equation, which possesses conserved quantities; in particular, $\lvert A(T)\rvert^{2}=\lvert A_{0}\rvert^{2}$, see Appendix C. It follows that, at the phase of the solution for which the momenta vanish ($\dot{y}_{n}=0$), Eq. (5) yields $4\epsilon^{2}\lvert A_{0}\rvert^{2}=\lVert\vec{y}\rVert^{2}/\lVert\vec{u}\rVert^{2}$ when retaining only the first order in $\epsilon$. Consequently,
@@
Lets us now perform numerical verifications of the theory above. We implement a pseudo-arclength solver [39, 40, 41] to numerically find families of exact nonlinear solutions of Eq. (1). The initial point for the pseudo–arclength solver is obtained using a shooting method, which searches for a small–amplitude exact solution with frequency $\Omega$, using the corresponding linear skin mode as an initial guess [40, 41]. Note that the resulting nonlinear modes are obtained with phases chosen such that their momenta satisfy $\dot{y}_{n}=0$. For the direct numerical simulations of the equations of motion, we use a Runge–Kutta scheme of order $8$ based on the Dormand–Prince method (DOP853853), with a time step $\Delta t=0.05$ and a one-step tolerance of $10^{-11}$ [42, 43, 44]. Hereafter, we consider $N=24$, $\alpha=1$, $\gamma=0.25$ and terminate the pseudo-arclength continuation at $\lVert\vec{y}\rVert^{2}=3$ so as to avoid intricate bifurcation structures that emerge at large nonlinearities, which are beyond the scope of this work.
Figure 4: Dependence of the nonlinear frequency sensitivity factor $\mathcal{S}=\partial\widetilde{\Omega}/\partial S$ against the lattice size NN with $\widetilde{\Omega}=(\Omega-\omega)/\omega$ and $S=\lVert\vec{y}\rVert^{2}$. We compute $\mathcal{S}$ for fixed values $\gamma=0.25$ (red squares), $\gamma=0.4$ (blue dots), $\gamma=0.55$ (cyan triangles), and $\gamma=0.7$ (purple stars). The derivatives are evaluated using a two-point finite difference at $S=0.01$ and 0 [45]. The straight solid, dashed, dotted, and dashed–dotted lines guide the eye to $\mathcal{S}\sim e^{-qN}$ with $q=0.06$, $0.07$, $0.10$, and $0.15$ (see text for details). Inset: Dependence of $\mathcal{S}$ on $\gamma$ for three lattice sizes, $N=24$, $N=32$, and $N=42$, shown as the blue, red, and cyan dotted symbols. We find that $\mathcal{S}(\gamma)\sim r^{-2}$, $r^{-2.5}$ and $r^{-3}$ for $N=24$, $N=32$, and $N=42$ respectively. In all cases, the analytical predictions show as solid curves overlapping with the numerical results, see also Fig. 7.
@@
We numerically compute the dependence of this frequency sensitivity factor, $\mathcal{S}$ [Eq. (10)] against the lattice size, $N$ in case of the hardening nonlinearitly with $g=1$. Note that the first derivative is evaluated using a two points finite difference formula in the interval $\lVert\vec{y}\rVert^{2}\in(0,0.01]$ [45]. Figure 4 shows the results of these computations for the family of nonlinear skin breathing modes with $j=1$ considering four different values of the nonreciprocity, $\gamma=0.25$, $\gamma=0.4$, $\gamma=0.55$, and $\gamma=0.7$, corresponding respectively to the red squares, blue dots, cyan triangles, and purple stars in Fig. 4. We find that for all values of $\gamma$, the $\mathcal{S}$ tends to asymptotically decay as the lattice size growth. Assuming the scaling $\mathcal{S}\sim e^{-qN}$ reveals that the exponent $q$ depends on $\gamma$, i.e., $q=q(\gamma)$. Interestingly, we find that $q$ increases with increasing $\gamma$. Specifically, for $\gamma=0.25,~0.4,~0.55$ and $0.7$, we obtain $q(\gamma)=0.06,~0.07,~0.1$ and $0.15$, respectively, as indicated by the black solid, dashed, dotted, and dashed–dotted lines in Fig. 4. The exponent $q$ is extracted by fitting the last five numerical data points.
Moreover, the inset of Fig. 4 depicts similar computations as above, but varying this time the nonreciprocal strength $\gamma$ at fixed lattice sizes, $N=24$, $N=32$ and $N=42$ respectively the red squares, blue diamonds and cyan triangles. Following similar arguments as above, we expect $\mathcal{S}\sim r^{p}$, with $p=p(N)$. By numerical fitting the sensitivity using the power law $\mathcal{S}=\sigma r^{-p}$, we find that $\mathcal{S}(\gamma)\sim r^{-2}$, $r^{-2.5}$, and $r^{-3}$ for $N=24$, $N=32$, and $N=42$, respectively [47]. The analytical predictions from Eq. (10) (solid curves) are superimposed on the numerical data and show excellent agreement, consistent with Fig. 7 in Appendix B